Write the first four terms of the sequence an=64(1/4)^n, for n=0, 1, 2, 3,
How do I finish problem?

Answer by venugopalramana(3286)

Write the first four terms of the sequence an=64(1/4)^n, for n=0, 1, 2, 3,
How do I finish problem?
PUT N=0,1,2,3 IN THE ABOVE..WE GET
FOR N=0................A0=64(1/4)^0=64
FOR N=1..........A1=64(1/4)^1=64/4=16
FOR N=2..........A2=64(1/4)^2=64/16=4
FOR N=3..........A3=64(1/4)^3=64/64=1